OpenStudy (zenmo):

A point in rectangular coordinates is given. (-4, -3). Convert it to polar coordinates. How to do this? Much Appreciated!

OpenStudy (dumbcow):

\[r^2 = x^2 + y^2\] \[\tan \theta = \frac{y}{x}\] Note that (-4,-3) is in 3rd quandrant so angle must be between 180 and 270

OpenStudy (zenmo):

I found that r=5 and tan beta sign= -3/-4 = 3/4. What do you do next with the tan beta sign?

OpenStudy (dumbcow):

you have to use inverse tangent on your calculator \[\theta = \tan^{-1} \frac{3}{4}\]

OpenStudy (zenmo):

I gotten 36.86 on my calculator, but the solution says 3.78?

OpenStudy (dumbcow):

ok well remember what i said about the angle being in the right quadrant tan is equal and positive in 1st and 3rd quandrants add 180 to 36.86 ----> 216.86 looks like thay want answer in radians .... convert 216.86 to radians to get 3.78

OpenStudy (zenmo):

Ok, I got it now. Thanks!

OpenStudy (dumbcow):

yw :)